For the circuit on the right, find the following quantities:


a. The circuit current.


b. The total resistance of the circuit.


c. The value of the unknown resistance, R.


d. The voltage-drop across all resistors in the circuit.


e. The power dissipated by all resistors.



100Ω


150Ω


E=25V


Rt= 800Ω



30Ω


R

1
Expert's answer
2022-03-22T15:46:43-0400

(1)

"I_t=\frac{V}{R_t}=\frac{25}{800}=0.03125A"

(2)

Net resistance


"R_{net}=800
\Omega"

(3)

"R_t=R_1+R_2+R_3+R_4"

"R_t=100+150+30+R"

"800=100+150+30+R"

"R=520\Omega"

(4)

"V_1=100\times0.03125=3.125\Omega"

"V_2=150\times0.03125=4.6875\Omega"

"V_3=30\times0.03125=0.9375\Omega"

"V_4=520\times0.03125=16.25\Omega"

(5) Power


"P_1=i_1V_1=3.125\times0.03126=0.0976W"

"P_2=i_2V_2=4.6875\times0.03126=0.1465W"

"P_3=i_3V_3=0.9375\times0.03126=0.02930W"

"P_4=i_4V_4=16.25\times0.03126=0.5079W"